40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 10 01:10:56 2018, gautam, Update, LSC, Some vertex locking restored PRMIcarrier.png
    Reply  Tue Aug 14 00:27:55 2018, gautam, Update, LSC, Locking prep 
       Reply  Tue Aug 14 02:01:12 2018, gautam, Update, LSC, DRMI locking - partial success DRMI.png
Message ID: 14162     Entry time: Tue Aug 14 02:01:12 2018     In reply to: 14160
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI locking - partial success 

After tweaking the AS55 demod phase, SRM alignment, triggering settings, I got a few brief DRMI locks in tonight, I'm calling it a success (though this isn't really robust yet). The main things to do now are:

  • turn on all the boosts on the LSC loops - today I only managed to trigger the PRCL boost filters successfully without blowing up the lock.
  • measure all 3 loops, tweak gain as necessary.
  • Run some sensing lines, tune the demod phase.
  • The SRCL triggering is strange to me - SRCL loop is currently triggered on POP22_I, but the 2f1 buildup in the symmetric side does not say anything about the linearity of the SRCL error signal? Or are we just hoping the SRM is in the correct place and engaging the servo? Anyway, this setting seems to work but perhaps once the locking is more robust the triggering can be fixed.
  • do a quick NB - I expect the main change to be that the AS55_Q dark noise contribution would have gone up on account on the reduced amount of light at this port.

I think the main IFO characterization remaining to be done to determine the status of the IFO post vent is to measure the losses of the arm cavities. IMO, we will need to certainly fix the clipping at ETMY before we attempt some serious locking.

Attachment 1: DRMI.png  223 kB  Uploaded Tue Aug 14 03:01:18 2018  | Hide | Hide all
DRMI.png
ELOG V3.1.3-